# Revisiting the Canonicalization for Fast and Accurate Crystal Tensor Property Prediction

### v3

@@ 3.3 Our Proposed Framework: GoeCTP
Crystal Graph Construction. To enable networks to handle such infinite crystal structures $(\mathbf{A},\mathbf{F},\mathbf{H})$, it is typically necessary to employ graph construction methods that represent the interactions between infinite crystal structures and atoms using finite graph data. Here, we use the crystal graph construction from Yan et al. [50, 51] to describe the structure and relationships within crystals. A detailed explanation of the crystal graph construction can be found in Appendix A.2. Upon construction, the edge features are denoted as $e_{ij}$, while the node features, representing atomic properties, are denoted as $\bm{a}_{i}$.
Node and Edge Feature Embedding. Building on previous work [47, 51, 50], node features $\bm{a}_{i}$ are embedded into a 92-dimensional CGCNN feature vector $\bm{f}_{i}$. Edge features $\mathbf{e}_{ij}$ are decomposed into their magnitude $||\mathbf{e}_{ij}||_{2}$ and a normalized direction vector $\hat{\mathbf{e}}_{ij}$. The magnitude is further mapped to a term similar to potential energy, $-c/||\mathbf{e}_{ij}||_{2}$, through the application of a radial basis function (RBF) kernel for encoding [21]. Subsequently, $\mathbf{e}_{ij}$ are embedded into feature vector $\bm{f}_{ij}^{e}$.
Prediction Module. Since the R&R module is responsible for preserving equivariance, any predictive network can serve as the Prediction module, such as those proposed by Yan et al. [50], Taniai et al. [42], Yan et al. [49], Lee et al. [20], Wang et al. [44], among others. For better performance, we select eComformer [50], which has demonstrated excellent performance in single-value property prediction tasks, as our Prediction module. A detailed explanation of eComFormer can be found in Appendix A.2. Once processed through the stacked layers of eComFormer, the node features are aggregated to generate the crystal’s global features as follows: $\bm{G}^{\mathrm{final}}=\frac{1}{n}\sum_{1\leq i\leq n}\bm{f}_{i}^{\mathrm{final}}$.
Reverse R&R Module. The primary function of the Reverse R&R module is to generate the equivariant tensor property predictions based on the crystal global features from the Prediction module. First, the Reverse R&R module transforms $\bm{G}^{\mathrm{final}}$ into a tensor output, as follows:
$\bm{\varepsilon}=f_{MLP}(\bm{G}^{\mathrm{final}})$ (5)
$\bm{\varepsilon}^{\mathrm{final}}=f_{rp}(\bm{\varepsilon},\mathbf{Q}),$ (6)
where $f_{MLP}(\cdot)$ represents a multilayer perceptron (MLP) and the operation reshaping dimension. Next, $f_{rp}(\cdot)$ utilizes the orthogonal matrix $\mathbf{Q}$ obtained from the R&R Module to convert the tensor output $\bm{\varepsilon}$ into its final equivariant form denoted as $\bm{\varepsilon}^{\mathrm{final}}$.
This conversion $f_{rp}(\cdot)$ for predicting the dielectric tensor can be expressed by:
$\bm{\varepsilon}^{\mathrm{final}}=\mathbf{Q}\bm{\varepsilon}\mathbf{Q}^{\top}.$ (7)
@@
4 Experiments
4.1 Experimental setup
Datasets. In this work, we evaluate the performance of GoeCTP on three key tensor property prediction tasks: the second-order dielectric tensor, the third-order piezoelectric tensor, and the fourth-order elastic tensor, respectively. Further details of the dataset and experimental setup can be found in Appendix A.10.
Baseline Methods. We selected several state-of-the-art methods in the field of crystal tensor property prediction, i.e. MEGNET [1, 31], EGTNN [53], and GMTNet [51], as baseline methods.
@@ 4.1 Experimental setup
4.2 Experimental Results
Predicting Dielectric Tensors. The performance of various models in predicting the dielectric tensor is summarized in Table 1. While GoeCTP and GMTNet show close results in terms of Fnorm and EwT 25%, GoeCTP demonstrates higher values for EwT 10% and EwT 5%, with an improvement of 5% in EwT 5% compared to GMTNet. This indicates that GoeCTP delivers higher-quality predictions compared to GMTNet. Furthermore, as GoeCTP is designed to be a flexible framework, we also evaluated its combination with other models, such as iComFormer [50] and CrystalFormer [42]. Detailed results of these combinations are provided in Appendix A.11. In addition to these experimental results, we further investigate the impact of different canonical forms on tensor property prediction in crystalline materials. Detailed analyses can be found in Appendix A.16.
MEGNET ETGNN GMTNet GoeCTP (Ours)
Fnorm $\downarrow$ 3.71 3.40 3.28 3.23
EwT 25% $\uparrow$ 75.8% 82.6% 83.3% 83.2%
EwT 10% $\uparrow$ 38.9% 49.1% 56.0% 56.8%
EwT 5% $\uparrow$ 18.0% 25.3% 30.5% 35.5%
Table 1: Comparison of performance metrics between MEGNET, ETGNN, GMTNet, and GoeCTP on the dielectric dataset.
Predicting Piezoelectric Tensors. The experimental results for the elastic tensor dataset are shown in Table 2. Similar to the experimental results on the dielectric tensors dataset, it achieves optimal results on all evaluation metrics. Further details on the results of GoeCTP’s combination with other models can be found in Appendix A.11.
MEGNET ETGNN GMTNet GoeCTP (Ours)
Fnorm $\downarrow$ 0.504 0.515 0.450 0.431
EwT 25% $\uparrow$ 39.3% 36.3% 45.5% 46.9%
EwT 10% $\uparrow$ 27.1% 20.5% 38.1% 42.9%
EwT 5% $\uparrow$ 11.4% 13.0% 34.3% 39.7%
Table 2: Comparison of performance metrics between MEGNET, ETGNN, GMTNet, and GoeCTP on the piezoelectric dataset.
Predicting Elastic Tensors. Experimental results on the elastic tensor dataset are presented in Table 3. GoeCTP demonstrated outstanding performance in predicting higher-order, complex tensors, outperforming all baseline methods across every evaluation metric. Notably, it achieved the lowest Fnorm of 107.11 and improved all EwT metrics by an average of 6% compared to GMTNet, emphasizing its general applicability in predicting diverse tensor properties in materials science. Further details on the results of GoeCTP’s combination with other models can be found in Appendix A.11.
MEGNET ETGNN GMTNet GoeCTP (Ours)
Fnorm $\downarrow$ 143.86 123.64 117.62 107.11
EwT 25% $\uparrow$ 23.6% 32.0% 36.0% 42.5%
EwT 10% $\uparrow$ 3.0% 3.8% 7.6% 15.3%
EwT 5% $\uparrow$ 0.5% 0.5% 2.0% 7.2%
Table 3: Comparison of performance metrics between MEGNET, ETGNN, GMTNet, and GoeCTP on the elastic dataset.
eCom. (ori. data) eCom. (aug. data) GoeCTP (Ours) (ori. data) GoeCTP (Ours) (aug. data)
@@ 4.2 Experimental Results
Total Time (s) $\downarrow$ 26.03 26.01 26.23 26.18
Piezoelectric: Fnorm $\downarrow$ 0.431 0.527 0.431 0.431
Total Time (s) $\downarrow$ 37.83 37.23 38.01 37.71
@@
Total Time (s) $\downarrow$ 83.26 83.02 90.10 89.60
@@
Table 4: Ablation study for verifying the $O(3)$ equivariance with dielectric, piezoelectric, and elastic dataset. After training GoeCTP, we extracted its Prediction module (i.e., eComFormer) for comparative testing on two different test sets
@@
Verifying the $O(3)$ Equivariance. To evaluate the effectiveness of GoeCTP, we conducted experiments to verify the $O(3)$ equivariance of tensor properties. Specifically, after training GoeCTP, we extracted its Prediction module (i.e., eComFormer) for comparative testing on two different test sets (original test set and augmented test set). All crystals in the original test set were adjusted to the canonical form, while the augmented test set was generated by applying arbitrary $O(3)$ group transformations to all crystals in the original test set. The method for generating the corresponding orthogonal matrices is from Heiberger [10]. We then evaluated both the Prediction module (eComFormer) and GoeCTP on these two datasets and compared the performance metrics.
The results on the three tensor datasets are shown in Table 4, GoeCTP performed equally well on the augmented test set as on the original test set in all tensor datasets, indicating that it maintains strong $O(3)$ equivariance for tensor properties. In contrast, the performance of the eComFormer method significantly declined on the augmented test set, with Fnorm decreasing by nearly 45% in dielectric tensor dataset, nearly 22% in piezoelectric tensor dataset, and nearly 29% in elastic tensor dataset, demonstrating that it does not meet the $O(3)$ equivariance requirements for tensor properties.
MEGNET ETGNN GMTNet GoeCTP (Ours)
Total Time (s) $\downarrow$ 663 1325 1611 616
Time/batch (s) $\downarrow$ 0.043 0.085 0.103 0.039
Table 5: Efficiency comparison on the dielectric dataset.
@@
Additionally, runtime comparisons between GoeCTP and eComFormer on the test set revealed that GoeCTP did not result in a significant increase in runtime. This demonstrates that when integrated into single-value property prediction networks, GoeCTP incurs almost no additional computational cost, enhancing its practicality while maintaining efficiency.
MEGNET ETGNN GMTNet GoeCTP (Ours)
Total Time (s) $\downarrow$ 943 1220 5771 938
Time/batch (s) $\downarrow$ 0.060 0.078 0.370 0.060
Table 6: Efficiency comparison on the piezoelectric dataset.
Efficiency. The results presented in Table 5, Table 6, and Table 7 illustrate the running times for GoeCTP and baseline methods. On the dielectric dataset, GoeCTP completed the entire training process with only 38.2% of the time spent compared to GMTNet. On the piezoelectric dataset, GoeCTP completed the entire training process with only 16.3% of the time spent compared to GMTNet. On the elastic dataset, GoeCTP required less than 7.0% of the time spent by GMTNet to complete the entire training process.
MEGNET ETGNN GMTNet GoeCTP (Ours)
Total Time (s) $\downarrow$ 2899 4448 $>$ 36000 2422
Time/batch (s) $\downarrow$ 0.036 0.056 $>$ 0.452 0.030
Table 7: Efficiency comparison on the elastic dataset.
To achieve $O(3)$ equivariance for tensor properties, GMTNet’s network architecture relies on irreducible representations and tensor operations, which considerably reduce computational efficiency, especially for higher-order tensor property (elastic tensor) prediction tasks where the time cost of GMTNet increases sharply. In contrast, GoeCTP requires no special architectural design for tensor property prediction tasks; it predicts tensor properties of varying orders using only a MLP at the network’s output, ensuring both efficiency and scalability across different tensor orders.
5 Concluding Remarks
@@
In this work, we revisit crystal tensor property prediction through the lens of canonicalization. Built upon the notion of canonicalization, we proposed a novel $O(3)$-equivariant framework GoeCTP for fast and accurate crystal tensor prediction. Benefiting from the canonicalization mechanism, GoeCTP serves as a plug-and-play module that can be integrated with any existing scalar property prediction network, enabling it to predict tensor properties with negligible additional computational cost. The proposed approach achieves state-of-the-art predictive performance and superior efficiency across a variety of widely used benchmark datasets for crystal tensor properties.
Except for the excel performance, our current GoeCTP has some limitations remain, which include (1) The performance of GoeCTP is inherently dependent on the performance of the Prediction module. (2) GoeCTP is specifically tailored for the prediction of tensor properties in crystalline materials and has not yet been adapted for other material types (See Appendix A.12 for more limitations details). In future study, we intend to improve performance by integrating prior knowledge related to the independent components of tensor properties across various crystal systems (see Appendix A.9). Further discussion on the utilization of space group constraints on tensor properties can be found in Appendix A.13.

### v4

@@ Introduction
Crystal Graph Preprocessing.
To enable neural networks to effectively process infinite crystal structures $(\mathbf{A},\mathbf{F},\mathbf{H})$, it is necessary to employ crystal graph construction and node and edge feature embedding techniques that represent atomic interactions within a finite graph representation. In this work, we adopt the crystal graph construction approach introduced by yancomplete; yanspace, where atoms are represented as nodes and interatomic interactions as edges. The node and edge features are subsequently embedded into vector representations $\boldsymbol{f}_{i}$ and $\boldsymbol{f}_{ij}^{e}$ to capture atomic and geometric information. A detailed description of the crystal graph construction procedure and the feature embedding scheme is provided in Appendix A.2.
Prediction Module.
Since the R&R module is responsible for preserving equivariance, any predictive network can serve as the Prediction module, such as those proposed by yancomplete; taniaicrystalformer; yan2022periodic; lee2024density; wang2024conformal; ito2025rethinking, among others. In this work, we select eComFormer (yancomplete) as the Prediction module to illustrate the overall workflow of our proposed method. A detailed explanation of eComFormer can be found in Appendix A.2. Once processed through the stacked layers of eComFormer, the node features are aggregated to generate the crystal’s global features as follows: $\boldsymbol{G}^{\mathrm{final}}=\frac{1}{n}\sum_{1\leq i\leq n}\boldsymbol{f}_{i}^{\mathrm{final}}$.
Reverse R&R Module.
The primary function of the Reverse R&R module is to generate the equivariant tensor property predictions based on the crystal global features from the Prediction module. First, the Reverse R&R module transforms $\boldsymbol{G}^{\mathrm{final}}$ into a tensor output, as follows:
$\boldsymbol{\varepsilon}=f_{MLP}(\boldsymbol{G}^{\mathrm{final}})$ (4)
$\boldsymbol{\varepsilon}^{\mathrm{final}}=f_{rp}(\boldsymbol{\varepsilon},\mathbf{Q}),$ (5)
where $f_{MLP}(\cdot)$ represents a multilayer perceptron (MLP) and the operation reshaping dimension. Next, $f_{rp}(\cdot)$ utilizes the orthogonal matrix $\mathbf{Q}$ obtained from the R&R Module to convert the tensor output $\boldsymbol{\varepsilon}$ into its final equivariant form denoted as $\boldsymbol{\varepsilon}^{\mathrm{final}}$. This conversion $f_{rp}(\cdot)$ for predicting the dielectric tensor can be expressed by:
$\boldsymbol{\varepsilon}^{\mathrm{final}}=\mathbf{Q}\boldsymbol{\varepsilon}\mathbf{Q}^{\top}.$ (6)
@@
Experiments
Experimental setup
Dataset Sample size Fnorm Mean Fnorm STD Unit
Dielectric 4713 14.7 18.2 Unitless
Piezoelectric 2701 0.79 4.03 $\text{C}/\text{m}^{2}$
Elastic 25110 306.4 238.4 GPa
Table 1: Dataset statistics.
Datasets.
In this work, we evaluate the performance of GoeCTP on three key tensor property prediction tasks: the second-order dielectric tensor, the third-order piezoelectric tensor, and the fourth-order elastic tensor, respectively. The datasets for dielectric tensor and piezoelectric tensor prediction are derived from the data curated by yanspace, which originates from the JARVIS-DFT database. Since nearly half of the crystal samples in the piezoelectric dataset have piezoelectric tensor labels that are entirely zero, training GoeCTP directly on this dataset can lead to severe overfitting. Therefore, we filtered out all crystal samples with zero-valued piezoelectric tensor labels from piezoelectric dataset. For the elastic tensor prediction task, the dataset is also obtained from the JARVIS-DFT database and is publicly accessible as dft_3d through the jarvis-tools package  [^2] https://pages.nist.gov/jarvis/databases/. The statistical details of the datasets are presented in Table 1. Further details of the dataset and experimental setup can be found in Appendix A.6.
Baseline Methods.
We selected several state-of-the-art approaches in the field of crystal tensor property prediction, including EGTNN (zhong2023general) and GMTNet (yanspace), as baseline methods. Moreover, since GoeCTP is designed as a flexible framework, we further evaluated its performance when combined with different backbone models. In addition to the eComFormer (yancomplete) used in Figure 2, we also incorporated the recently proposed CrystalFramer (ito2025rethinking).
Evaluation Metrics.
@@
Experimental Results
Predicting Dielectric Tensors.
The performance of various models in predicting the dielectric tensor is summarized in Table 2. Since GoeCTP (C. Fra.) employs a more advanced backbone, its performance is notably superior to that of GoeCTP (C. eCom.). Both GoeCTP (eCom.) and GoeCTP (C. Fra.) outperform the baseline methods, demonstrating the effectiveness of the proposed GoeCTP framework. Comprehensive results combining GoeCTP with additional backbone architectures are provided in Appendix A.7. In addition to these experimental results, we further investigate the impact of different canonical forms on tensor property prediction in crystalline materials. Detailed analyses can be found in Appendix A.11.
ETGNN GMTNet GoeCTP (eCom.) GoeCTP (C. Fra.)
Fnorm $\downarrow$ 3.40 3.28 3.23 3.05
EwT 25% $\uparrow$ 82.6% 83.3% 83.2% 86.4%
EwT 10% $\uparrow$ 49.1% 56.0% 56.8% 62.6%
EwT 5% $\uparrow$ 25.3% 30.5% 35.5% 43.5%
Table 2: Comparison on the dielectric dataset. Lower Fnorm and higher EwT indicate better performance. GoeCTP (eCom.) denotes GoeCTP using eComFormer as the backbone, while GoeCTP (C. Fra.) refers to GoeCTP with CrystalFramer as the backbone.
Predicting Piezoelectric Tensors.
The experimental results for the piezoelectric tensor dataset are shown in Table 3.
Compared to the results on the dielectric dataset, predicting the piezoelectric tensor presents a greater challenge. As shown in Table 1, the overall mean Fnorm of the piezoelectric tensor is relatively small, which leads to lower overall EwT values. Consequently, achieving higher EwT requires more accurate predictions from the model. GoeCTP (C. Fra.) achieves relatively high EwT at both 5% and 10%, indicating that it attains high prediction accuracy for a subset of samples. Meanwhile, GMTNet and GoeCTP (eCom.) obtain better Fnorm results, suggesting that their overall prediction errors are comparatively smaller.
ETGNN GMTNet GoeCTP (eCom.) GoeCTP (C. Fra.)
Fnorm $\downarrow$ 0.873 0.752 0.778 0.827
EwT 25% $\uparrow$ 0% 6.29% 2.59% 3.33%
EwT 10% $\uparrow$ 0% 1.48% 1.14% 2.59%
EwT 5% $\uparrow$ 0% 1.11% 0.04% 2.22%
Table 3: Comparison on the piezoelectric dataset.
Predicting Elastic Tensors.
Experimental results on the elastic tensor dataset are summarized in Table 4. Consistent with the observations on the dielectric dataset, both GoeCTP (eCom.) and GoeCTP (C. Fra.) achieved outstanding performance in predicting higher-order, complex tensors, surpassing all baseline methods across all evaluation metrics. Notably, GoeCTP (C. Fra.) delivered a further improvement over GoeCTP (eCom.). Additional results demonstrating GoeCTP combined with other backbone models are provided in Appendix A.7.
ETGNN GMTNet GoeCTP (eCom.) GoeCTP (C. Fra.)
Fnorm $\downarrow$ 123.64 117.62 107.11 95.98
EwT 25% $\uparrow$ 32.0% 36.0% 42.5% 49.7%
EwT 10% $\uparrow$ 3.8% 7.6% 15.3% 20.3%
EwT 5% $\uparrow$ 0.5% 2.0% 7.2% 9.8%
Table 4: Comparison on the elastic dataset.
eCom. (ori.) eCom. (aug.) GoeCTP (ori.) GoeCTP (aug.)
@@
Dielectric: Test Time (s) $\downarrow$ 26.03 26.01 26.23 26.18
Piezoelectric: Fnorm $\downarrow$ 0.778 0.957 0.778 0.778
Piezoelectric: Test Time (s) $\downarrow$ 5.21 5.40 5.53 6.05
@@
Elastic: Test Time (s) $\downarrow$ 83.26 83.02 90.10 89.60
@@
Table 5: Ablation study for verifying the $O(3)$-equivariance using dielectric, piezoelectric, and elastic datasets. After training GoeCTP (eCom.), we extracted its prediction module (i.e., eComFormer) for comparative testing on two different test sets.
Verifying the $O(3)$ Equivariance.
@@
To evaluate the effectiveness of GoeCTP, we conducted experiments to verify the $O(3)$ equivariance of tensor properties. Specifically, after training GoeCTP (eCom.), we extracted its Prediction module (i.e., eComFormer) for comparative testing on two different test sets (original test set and augmented test set). All crystals in the original test set were adjusted to the canonical form, while the augmented test set was generated by applying arbitrary $O(3)$ group transformations to all crystals in the original test set. The method for generating the corresponding orthogonal matrices is from heiberger1978algorithm. We then evaluated both the Prediction module (eComFormer) and GoeCTP (eCom.) on these two datasets and compared the performance metrics.
@@
The results on the three tensor datasets are shown in Table 5, GoeCTP performed equally well on the augmented test set as on the original test set in all tensor datasets, indicating that it maintains strong $O(3)$ equivariance for tensor properties. In contrast, the performance of the eComFormer method significantly declined on the augmented test set, demonstrating that it does not meet the $O(3)$ equivariance requirements for tensor properties. Additionally, runtime comparisons between GoeCTP and eComFormer on the test set revealed that GoeCTP did not result in a significant increase in runtime. This demonstrates that when integrated into scalar property prediction networks, GoeCTP incurs almost no additional computational cost, enhancing its practicality while maintaining efficiency.
Dielectric: ETGNN GMTNet GoeCTP (eCom.) GoeCTP (C. Fra.)
Total Time (s) $\downarrow$ 1325 1611 616 1976
Time/epoch (s) $\downarrow$ 6.625 8.055 3.08 9.88
Piezoelectric: ETGNN GMTNet GoeCTP (eCom.) GoeCTP (C. Fra.)
Total Time (s) $\downarrow$ 736 7012 356 5768
Time/epoch (s) $\downarrow$ 2.18 35.06 1.78 5.77
Elastic: ETGNN GMTNet GoeCTP (eCom.) GoeCTP (C. Fra.)
Total Time (s) $\downarrow$ 4448 >> 36000 2422 17197
Time/epoch (s) $\downarrow$ 22.24 >> 180 12.11 34.394
Table 6: Efficiency comparison.
Efficiency.
Table 6 summarizes the training efficiency of GoeCTP and baseline methods. On the dielectric, piezoelectric, and elastic datasets, GoeCTP (eCom.) required only 38.2%, 5%, and 7% of GMTNet’s training time, respectively. Although GoeCTP (C. Fra.) achieved higher accuracy at the cost of longer training, it still maintained a efficiency advantage over GMTNet, particularly on high-order tensor datasets such as the elastic dataset.
To achieve $O(3)$ equivariance for tensor properties, GMTNet employs a network architecture based on irreducible representations and tensor operations. Moreover, its final tensor output is obtained through gradient computation, which may significantly reduce computational efficiency. This inefficiency becomes particularly pronounced when computing gradients for high-order tensors such as elastic tensors. Beyond training costs, the gradient computation also may lead to slower inference. For instance, in the elastic tensor task, GMTNet requires more than 0.478 seconds per material, whereas GoeCTP (eCom.) requires only 0.036 seconds per material. In contrast, GoeCTP achieves equivariant tensor predictions without any specialized architectural constraints. It utilizes a simple multilayer perceptron (MLP) at the network output to predict tensor properties of various orders, ensuring high efficiency and scalability across different tensor orders.
Conclusion
@@
In this work, we revisit crystal tensor property prediction through the lens of canonicalization. Built upon the notion of canonicalization, we proposed a novel $O(3)$-equivariant framework GoeCTP for fast and accurate crystal tensor prediction. Benefiting from the canonicalization mechanism, GoeCTP serves as a plug-and-play module that can be integrated with any existing scalar property prediction network, enabling it to predict tensor properties with negligible additional computational cost.
Except for the excel performance, our current GoeCTP has some limitations remain:
  1. 1.
    GoeCTP is a plug-and-play $O(3)$-equivariant framework designed to enhance the backbone network’s capability for equivariant prediction. Consequently, its overall performance inherently depends on the strength of the chosen backbone. For example, GoeCTP (eCom.) demonstrates higher efficiency, while GoeCTP (C. Fra.) achieves greater predictive accuracy. Striking an optimal balance between efficiency and accuracy remains a topic for further investigation.
  2. 2.
    As demonstrated in higham1986computing and Theorem 1, polar decomposition applied to an invertible matrix $\mathbf{L}$ produces a unique canonical form $\mathbf{H}$. For 3D crystal structures, the lattice matrix $\mathbf{L}$ is always full rank (i.e., invertible), ensuring the applicability of our approach to 3D crystal systems. However, in special cases such as 2D single-layer crystals (novoselov2005two; sherrell2022bright), the lattice matrix $\mathbf{L}$ may be rank-deficient. In these scenarios, polar decomposition may fail to yield a unique canonical form $\mathbf{H}$, making it impossible to consistently align crystals with different spatial orientations to a common canonical form, thereby hindering $O(3)$-equivariance.
  3. 3.
    The current framework underutilizes prior knowledge related to space group constraints and does not yet ensure that predicted tensor results strictly adhere to space group symmetry constraints. In future work, we plan to improve model performance by integrating priors concerning the independent tensor components across different crystal systems (see Appendix A.5). Further discussion on incorporating space group symmetry constraints into tensor property prediction is provided in Appendix A.8.
Acknowledgments
This work was partially supported by the Research Grants Counil (RGC) of the Hong Kong (HK) SAR (Grant No. 15208725 and 15208222), the Young Scientists Fund of National Natural Science Foundation of China (NSFC) (Grant No. 62206235), and the Hong Kong Polytechnic University (Grant No. A0046682 and P0057774).